OpenStudy (anonymous):

help please trying to finish Match each set of ordered pairs with their rise over run. Example: 1=A Match Term Definition 1: (4,-2), (3,1) A) -3 2: (4,-2), (-3,1) B) -1 3: (-4,2), (-3,1) C) -1/7 4: (-4,2), (3,1) D) -3/7

OpenStudy (anonymous):

1 and A, 2 and D, 3 and C, 4 and B

OpenStudy (triciaal):

rise = change in y = y2 - y1 run = change in x = x2 - x1

OpenStudy (anonymous):

Thank you both

OpenStudy (triciaal):

welcome